OpenStudy (anonymous):

put the equation in the form of ax^2+bx+c=0 x^2+8x+18=3

OpenStudy (anonymous):

Seriously? This is easy. C'mon! Now your not even trying :/

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

PUT IT IN THE FORM. Don't tell me what you subtract by

OpenStudy (anonymous):

x^2+8x+15=0

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

wait i did the equation wrong i mean i wrote it wrong xD

OpenStudy (anonymous):

WHAT!?!?!

OpenStudy (anonymous):

you did THIS equation right

OpenStudy (anonymous):

x^2+8x+10=3 -3 x^2+8x+7=0 this is the right equation xD

OpenStudy (anonymous):

jesus christ. WHAT DID I TELL YOU! I'm probs acting like your mother now. BUT NO RUSHING

OpenStudy (anonymous):

and yes, correct :P

OpenStudy (anonymous):

yup :p